Question 4 20 pts Stock XYZ has a current dividend of $3.00 The dividend is expected to grow forever at a rate of 3.00% Basedplease answer
Answer #1

D2 = 3(1.03)2 = $3.183

Calculating Stock Price in Year 1 using Constant Growth Model,

Stock Price in Year 1 = D2/(r - g) = 3.183/(0.04 - 0.03)

Stock Price in Year 1 = $318.30

Dividend Yield in Year 1 = 3.183/318.30

Dividend Yield in Year 1 = 0.01

Dividend Yield in Year 1 = 1.00%